How do I go about setting up this equation?
2 trains are a mile and a 1/2 apart
A is traveling at 3/4ths the speed of train B
they'll collide in 4 1/2 minutes
how fast is each train going
Thanks in advance.

Answer by Alan3354(69443)

How do I go about setting up this equation?
2 trains are a mile and a 1/2 apart
A is traveling at 3/4ths the speed of train B
they'll collide in 4 1/2 minutes
how fast is each train going
----------------------
The fact that they'll collide implies they're moving toward each other, so the speeds are added.
---
A = 0.75B
A + B = 1.75B (the "closing speed")
------
d = r*t
1.5 = 1.75B*4.5
B = 1.5/(1.75*4.5) = 1/5.25 miles/minute
-----
A = 0.75B = 1/7 miles/minute
==================
Multiply by 60 to get miles per hour